### v1

@@ Abstract
The ability of robots to manipulate objects relies heavily on their aptitude for visual perception. In domains characterized by cluttered scenes and high object variability, most methods call for vast labeled datasets, laboriously hand-annotated, with the aim of training capable models. Once deployed, the challenge of generalizing to unfamiliar objects implies that the model must evolve alongside its domain. To address this, we propose a novel framework that combines Semi-Supervised Learning (SSL) with Learning Through Interaction (LTI), allowing a model to learn by observing scene alterations and leverage visual consistency despite temporal gaps without requiring curated data of interaction sequences. As a result, our approach exploits partially annotated data through self-supervision and incorporates temporal context using pseudo-sequences generated from unlabeled still images. We validate our method on two common benchmarks, ARMBench [34] mix-object-tote and OCID [37], where it achieves state-of-the-art performance. Notably, on ARMBench, we attain an $\text{AP}_{50}$ of $86.37$, almost a $20\%$ improvement over existing work, and obtain remarkable results in scenarios with extremely low annotation, achieving an $\text{AP}_{50}$ score of $84.89$ with just $1\%$ of annotated data compared to $72$ presented in [34] on the fully annotated counterpart  [^1] Code available in https://github.com/mkimhi/RISE .
@@ 1 Introduction
Figure 1: Method: Pseudo-sequence generation from a single unlabeled image. The input is weakly augmented to produce the “before” image $x_{1}$ and augmented again to yield the “after” image $x_{2}$. To emulate scene interaction, objects are drawn from the object memory bank, transformed, and inserted into the “after” frame. The segmentation model’s task is to simultaneously associate objects that persist between frames (subject to occlusion), maintain consistency of object instance embedding, and correctly predict the ground-truth mask of the added objects.
@@
The resulting framework can be considered the first to employ self-supervised learning through interaction, achieving better performance than each paradigm individually. We set a new state-of-the-art on the ARMBench [34] benchmark and OCID [37] (RGB only). Notably, our method trained on $1\%$ of annotated data surpasses the performance of the well-established Deformable DETR [53] architecture, even when trained on 10×10$\times$10 × additional annotated data (improving $+16.86$ AP using Swin-L Transformer as feature extractor).
@@ 3.2 Learning Through Interaction
where $(u,v)$ is the top-left corner where the object is inserted, drawn from distribution $Beta(\alpha,\beta)\in\mathbb{R}^{2}$, and $w$, $h$ are the feasible horizontal and vertical regions that ensure that the object is contained within the frame (see Appendix A.A). The choice of $Beta$ and its parameters reduces the likelihood of objects inserted near the center, where they might obstruct unlabeled objects. Another safeguard prevents inserting an object if it would overlap with any of the previously inserted objects by more than $85\%$.
@@ 3.3 Self-Supervision
In Appendix A.D, we demonstrate that early in training, setting the quantile too low lets in more predictions of low-quality signals, interfering with the model. Alternatively, setting the bar (too) high [36] risks missing out on meaningful supervision signals.
@@
The threshold $\gamma_{t}^{cls}$ discards instances with class scores $\hat{c}_{i}$ below it and tightens over time (training steps $t$). Conversely, the quantile $Q(p_{t})$ loosens over time with its probability $p_{t}=0.995\cdot(1-\nicefrac{{t}}{{T}})$ decays over subsequent training steps $t$, with $T$ denoting the total number of training steps. As a result, the quantile allows more predictions into the model as training progresses. This strategy can mitigate incorrect model beliefs and reduce confirmation biases. We evaluate this strategy quantitatively and demonstrate its advantage over thresholds and quantiles in Tab. 4, with additional details in Appendix A.C. We recognize that exploring different quantile strategies may further improve self-supervision and set it aside for future work.
@@ 3.4 Unified Framework
where $\mathds{1}$ indicates that the supervised loss $\mathcal{L}_{\mathrm{s}}$ and unsupervised loss $\mathcal{L}_{\mathrm{u}}$ are used according to the availability of ground-truth labels $\mathbf{y}$, and $\mathcal{L}_{embed}$ denotes the weighted combinations of the association loss. Hyperparameter search for $\lambda_{3}$ and details on $\lambda_{1},\lambda_{2}$ and $\lambda_{4}$ are provided in Appendix A.A.
@@ 4.1 Setup
The OCID [37] dataset (containing $2,\!390$ images and $31$ classes) for various rates of labeled-to-unlabeled data, and compare it to the current state-of-the-art[32].
@@
4.2 Results
Tab. 1 shows the results of RISE on various data partitions of labeled/unlabeled ratios of the ARMBench, compared with Deformable DETR and SAM[19] (fine-tuned), as well as the results reported by the authors of ARMBench [34]. Both DeDETR and RISE use Swin-L[28] ($197$M parameters) as backbone, while SAM uses ViT-H[23] ($636$M parameters) and RoboLLM[29] uses Beit-3 base (87M parameters). Across all partitions, RISE outperforms the other methods. Fig. 4 illustrates high-quality masks predicted by RISE trained on $1\%$ of the labeled, with $99\%$ of the remaining data treated as unlabeled. Most of the line-of-sight objects are accurately segmented and a few heavily occluded objects are missed. Importantly, RISE trained on $1\%$ annotated data performs better than DeDETR and SAM trained on $10\%$ annotated samples (10×10$\times$10 × the amount of annotations for training/fine-tuning).
@@ 5 Conclusion
In this work, we present RISE, a novel framework that incorporates semi-supervised learning with learning through scene interaction in the context of a few-annotation data regime. RISE is modular and can complement other segmentation models that emit intermediate instance embedding. We demonstrate that RISE improves $\text{AP}_{50}$ by over $+10$ compared to previous state-of-the-art, after training end-to-end on just $0.5\%$ of the labeled data (with $99.5\%$ of the data treated as unlabeled). With just $1\%$ of the labeled data, RISE achieves better performance than the baselines (DeDETR, SAM, RoboLLM) trained on 10×10$\times$10 × the amount of labeled data. On OCID (RGB), RISE sets a new state-of-the-art, and is near state-of-the-art when restricted to a fraction of the annotations. For future work, we intend on leveraging “before” and “after” observations directly using robotic item grasping in real-world environments (rather than synthetically inserting instances into images), with the overarching goal of lifelong learning for robot perception.
@@
Appendix A.A Technical details
@@
Appendix A.B Prediction Matching
@@ Appendix A.C Thresholds
Appendix A.D Study of Cascade Threshold
@@
Appendix A.E Failure Cases

### v2

@@ Abstract
The ability of robots to manipulate objects relies heavily on their aptitude for visual perception. In domains characterized by cluttered scenes and high object variability such as traffic, navigation and object grasping, most methods call for vast labeled datasets, laboriously hand-annotated, with the aim of training capable models. Once deployed, the challenge of generalizing to unfamiliar objects implies that the model must evolve alongside its domain. To address this, we propose a novel framework that combines Semi-Supervised Learning (SSL) with Learning Through Interaction (LTI), allowing a model to learn by observing scene alterations and leverage visual consistency despite temporal gaps without requiring curated data of interaction sequences. As a result, our approach exploits partially annotated data through self-supervision and incorporates temporal context using pseudo-sequences generated from unlabeled still images. We validate our method on two common benchmarks, ARMBench mix-object-tote and OCID, where it achieves state-of-the-art performance. Notably, on ARMBench, we attain an $\text{AP}_{50}$ of $86.37$, almost a $20\%$ improvement over existing work, and obtain remarkable results in scenarios with extremely low annotation, achieving an $\text{AP}_{50}$ score of $84.89$ with just $1\%$ of annotated data compared to previous state of the art of $82$ which targeted the fully annotated dataset.
@@ 1 Introduction
Figure 1: Pseudo-sequence generation from a single unlabeled image. The input is weakly augmented to produce the “before” and “after” images $x_{1}$ and $x_{2}$. To emulate scene interaction, objects are drawn from the object memory bank, transformed, and inserted into the “after” frame. The segmentation model’s task is to simultaneously associate objects that persist between frames (subject to occlusion), maintain consistency of object instance embedding, and correctly predict the ground-truth mask of the added objects.
@@
The resulting framework can be considered the first to employ self-supervised learning through interaction, achieving better performance than each paradigm individually. We set a new state of the art on the ARMBench [35] benchmark and OCID [38] (RGB only). Notably, our method trained on $1\%$ of annotated data surpasses the performance of the well-established Deformable DETR [54] architecture, even when trained on 10×10$\times$10 × additional annotated data (improving $+16.86$ AP using Swin-L Transformer as feature extractor).
@@ 3.2 Learning Through Interaction
where $(u,v)$ is the top-left corner where the object is inserted, drawn from distribution $Beta(\alpha,\beta)\in\mathbb{R}^{2}$, and $w$, $h$ are the feasible horizontal and vertical regions that ensure that the object is contained within the frame (see Appendix A). The choice of $Beta$ and its parameters reduces the likelihood of objects inserted near the center, where they might obstruct unlabeled objects. Another safeguard prevents inserting an object if it would overlap with any of the previously inserted objects by more than $85\%$.
@@ 3.3 Self-Supervision
In Appendix E, we demonstrate that early in training, setting the quantile too low lets in more predictions of low-quality signals, interfering with the model. Alternatively, setting the bar (too) high [37] risks missing out on meaningful supervision signals.
@@
The threshold $\gamma_{t}^{cls}$ discards instances with class scores $\hat{c}_{i}$ below it and tightens over time (training steps $t$). Conversely, the quantile $Q(p_{t})$ loosens over time with its probability decays over subsequent training steps $t$, with $T$ denoting the total number of training steps. As a result, the quantile allows more predictions into the model as training progresses. This strategy can mitigate incorrect model beliefs and reduce confirmation biases. We evaluate this strategy quantitatively and demonstrate its advantage over thresholds and quantiles in Tab. 4, with additional details in Appendix C. We recognize that exploring different quantile strategies may further improve self-supervision and set it aside for future work.
@@ 3.4 Unified Framework
where $\mathds{1}$ indicates that the supervised loss $\mathcal{L}_{\mathrm{s}}$ and unsupervised loss $\mathcal{L}_{\mathrm{u}}$ are used according to the availability of ground-truth labels $\mathbf{y}$, and $\mathcal{L}_{embed}$ denotes the weighted combinations of the association loss. Hyperparameter search for $\lambda_{3}$ and details on $\lambda_{1},\lambda_{2}$ and $\lambda_{4}$ are provided in Appendix A.
@@ 4.1 Setup
The OCID [38] dataset (containing $2,\!390$ images and $31$ classes) for various rates of labeled-to-unlabeled data, and compare it to the current state of the art±[32].
@@
4.2 Results
Tab. 1 shows the results of RISE on various data partitions of labeled/unlabeled ratios of the ARMBench, compared with Deformable DETR and SAM[20] (fine-tuned), as well as the results reported by the authors of ARMBench [35]. Both DeDETR and RISE use Swin-L[29] ($197$M parameters) as backbone, while SAM uses ViT-H[24] ($636$M parameters) and RoboLLM[30] uses Beit-3 base (87M parameters). Across all partitions, RISE outperforms the other methods. Fig. 3 illustrates high-quality masks predicted by RISE trained on $1\%$ of the labeled, with $99\%$ of the remaining data treated as unlabeled. Most of the line-of-sight objects are accurately segmented and a few heavily occluded objects are missed. Importantly, RISE trained on $1\%$ annotated data performs better than DeDETR and SAM trained on $10\%$ annotated samples (10×10$\times$10 × the amount of annotations for training/fine-tuning). As an additional baseline we compare RISE with the “Segment Anything” (SAM) [20], for more details refer to Appendix D
@@ 5 Conclusion
In this work, we present RISE, a novel framework that incorporates semi-supervised learning with learning through scene interaction in the context of a few-annotation data regime. RISE is modular and can complement other segmentation models that emit intermediate instance embedding. We demonstrate that RISE improves $\text{AP}_{50}$ by over $+10$ compared to previous state of the art, after training end-to-end on just $0.5\%$ of the labeled data (with $99.5\%$ of the data treated as unlabeled). With just $1\%$ of the labeled data, RISE achieves better performance than the baselines (DeDETR, SAM, RoboLLM) trained on 10×10$\times$10 × the amount of labeled data. On OCID (RGB), RISE sets a new state of the art, and is near state of the art when restricted to a fraction of the annotations. For future work, we intend on leveraging “before” and “after” observations directly using robotic item grasping in real-world environments (rather than synthetically inserting instances into images), with the overarching goal of lifelong learning for robot perception.
@@
Appendix A Technical details
@@
Appendix B Prediction Matching
@@ Appendix C Thresholds
Appendix D Foundation Model Comparison
@@
Appendix E Study of Cascade Threshold
Appendix F Failure Cases
@@
Acknowledgment: Bin illustrations in figure Fig. 2 are attributed and modified from [1].